# Reconstructing Materials Tetrahedron: Challenges in Materials Information Extraction

### v2

@@ Abstract
Discovery of new materials has a documented history of propelling human progress for centuries and more. The behaviour of a material is a function of its composition, structure, and properties, which further depend on its processing and testing conditions. Recent developments in deep learning and natural language processing have enabled information extraction at scale from published literature such as peer-reviewed publications, books, and patents. However, this information is spread in multiple formats, such as tables, text, and images, and with little or no uniformity in reporting style giving rise to several machine learning challenges. Here, we discuss, quantify, and document these challenges in automated information extraction (IE) from materials science literature towards the creation of a large materials science knowledge base. Specifically, we focus on IE from text and tables and outline several challenges with examples. We hope the present work inspires researchers to address the challenges in a coherent fashion, providing a fillip to IE towards a materials knowledge base.
@@ 1 Introduction
Understanding a material’s behaviour requires knowledge about its composition, properties, processing and testing protocols, and microstructure—represented as the materials science (MatSci) tetrahedron (see Fig. 1). These different aspects of a material are reported by researchers in peer-reviewed publications, patents, and other scientific documents. Recently, there have several attempts to exploit the advances in machine learning (ML) and artificial intelligence (AI) towards automated information extraction (IE) from literatureKononova et al. (2021); Zaki et al. (2023a); Ravinder et al. (2021); Zaki et al. (2024). These include the development of materials specific language models Gupta et al. (2022); Huang and Cole (2022); Trewartha et al. (2022); Gupta et al. (2023), rule-based systems Swain and Cole (2016); Mavracic et al. (2021); Zaki et al. (2022b); Venugopal et al. (2021); Zaki et al. (2022c), IE from tables Gupta et al. (2023); Zhao and Cole (2022); Zhao et al. (2023), and IE from images Mukaddem et al. (2019); Zhang and Shao (2022); Zaki et al. (2023b, 2022a). The widely varying information expression styles in research papers makes the automated MatSci IE a challenging task. Most of the works have been focused on IE in a specific domain and hence the transferability to different materials is not explored. Moreover, there exists no consolidated work that explores the specific challenges associated with IE in MatSci and the gain associated with solving these challenges, which provides a clear direction to the researchers regarding the areas that require increased attention.
@@
We thoroughly review MatSci articles to identify IE challenges towards completing the materials tetrahedron (see Fig. 1). We also highlight some of the major challenges toward the development of a “universal” MatSci knowledge base linking the extracted information from multiple sources and forms of data—structured, semi-structured, and unstructured. Indeed, there exist millions of scientific documents reporting information about various materials known to humans. Thus, the automated development of MatSci IE will lead to a rich knowledge base on materials. The outline of the paper is as follows: First, we explain the methodology of collecting papers for review and annotation process. Then, in the results and discussion sections, we investigate the proportion of each of the entities, such as composition, structure, properties, processing and testing conditions, reported in tables or text of the articles, followed by the challenges faced in their extraction. We quantify how frequently a challenge occurs to motivate researchers to gauge the amount of information that will be obtained after solving the respective challenges. We further identify the challenges in extracting and connecting the information from text and tables, and also among different tables belonging to the same MatSci research papers. Finally, we provide some guidelines for presenting machine and human-friendly tables that enable automated MatSci IE from research papers.
@@ 2 Methodology
To elucidate the challenges, we referred to a dataset of 2536 peer-reviewed publications on MatSci. This dataset is based on recent work on IE from tables Gupta et al. (2023), where the training data was annotated using distant supervision while the val and test data were annotated manually. Fig. 2 shows different sections of the paper where these different components are majorly reported. The statistics of each challenge were computed by randomly taking 50/100 tables from the manually annotated val and test dataset. In cases, where this was not applicable, we further performed manual annotation on additional 50 papers or 100 relevant tables, selected randomly from the corpus. For instance, for composition extraction, we randomly selected 100 composition tables from the manual annotation in the existing dataset. However, no such manual annotation was available for properties. For this problem, we selected 100 random property tables from the corpus and manually annotated the frequency of the challenges in property extraction. Note that all the challenges are based on manual annotation and hence are reliable than any ML-based technique such as distant supervision. Further, we manually analyzed tables or text for the occurrence of each of the entities such as composition, structure, and property. All the results and data associated with the annotation process are shared in the following link.
@@ 3.1.1 Extracting compositions from tables
Here, we summarize the major challenges in IE from tables. To this extent, we investigated 100 randomly selected composition tables from the manually annotated data to report the frequency of occurrence of each challenge.
@@
a. Variation in table structure and information content: An analysis of 100 random MatSci composition tables revealed that these tables do not follow any standard structure. Accordingly, following earlier schema proposed by Gupta et al. Gupta et al. (2023), composition tables can categorized into two broad categories—multi-cell composition (MCC) and single-cell composition (SCC) tables. These are further subdivided into tables containing complete information (CI) and partial information (PI). When the entire composition is written inside a single cell, it is classified as an SCC table, whereas when the composition is written across multiple cells of the table by reporting the value of each constituent (compounds or elements) of the composition in separate cells, it is defined to be an MCC table. If the table contains all the information regarding the constituents of the material, they are classified as CI tables (complete information). Alternatively, if only some of the constituents are mentioned in the table for the material, they are PI tables. In the latter case, we need to extend the analysis to the text of the article to extract the full composition. Fig. 4 illustrates all 4 types of tables Nazabal et al. (2003); Zaharescu et al. (2008); Narayanan and Shashikala (2015); Jestin et al. (2003). The most prevalent composition table types are MCC-CI (36%), followed by SCC-CI (30%). PI tables are less common, with 24% being present as MCC-PI and the remaining 10% as SCC-PI. Note that this distribution may vary significantly depending on the material types as well. For instance in alloys, it is a common practice to skip the major element while describing the composition in a table. In previous work by Gupta et al.. Gupta et al. (2023), while an F1 score of 78.21% and 65.41% have been achieve for extraction from SCC-CI and MCC-CI tables, respectively, an F1 score of only 51.66% has been achieved for extraction from MCC-PI. Extraction from SCC-PI is an open challenge.
@@
b. Presence of nominal and experimental compositions: While the nominal composition is the amount of chemicals taken initially to prepare the material, analyzed/experimental composition refers to the actual composition of material obtained after analyzing the manufactured material (see Fig. 5(a)) Youngman et al. (2000); Petkov and Ewen (1999). Our analysis revealed that in 3% of the tables, both nominal and analyzed/experimental compositions are reported. These values are not reported in any fixed pattern, making it difficult to separate the nominal and analyzed compositions correctly after extraction.
c. Composition in table headers: Most tabular IE models like Tabbie Iida et al. (2021), DiSCoMaT Gupta et al. (2023) perform better when row/column headers contain appropriate information regarding its contents. In MatSci tables, the headers are mostly material IDs, compound names, properties, processing and testing labels, and the inner cells contain corresponding values. However, in 6% of the tables, we found that the compounds with their values were present in the heading, which makes it hard for the IE models to extract the desired information. For instance, SeGePb Jain et al. (2009) or = 10%, = 20%,… Minati et al. (2007) are column headers which contain both the compounds and corresponding concentration in the heading. 67% of these were SCC-CI, whereas the rest 33% were MCC-PI tables.
@@
d. Compositions and related info inferred from other documents: In some tables, the details of the compositions studied are not explicitly mentioned; rather references to previous research publications which uses the same material is provided in the tables or their captions (see Fig. 5(b)). Thus, such a task requires the composition to be extracted from a previous work, which then should be combined with the property in the present work. Such references were seen in 11% of the tables Brasil and Meyer (2004); Chen et al. (2003).
e. Compounds inferred from material ID: 10% of the tables have IDs in which valuable information regarding the composition was present (see Fig. 6. For example, some of the material IDs have the molar ratios of compounds present in the material, while others have the composition written in an abbreviated form Nazabal et al. (2001). In such cases, compositions are not explicitly mentioned in the table. There are also tables where the composition of the materials is not specified; instead, their standard names are used as IDs. Such examples include Wollastonite or Diopside Schneider et al. (2000), which has a fixed chemical composition and therefore the ID can be used to infer the composition of the material.
@@
f. Variables used to represent compounds: When a composition is expressed with variables such as TeO+15BO+15PO+LiO, where = 5, 10, 15, 20, 25 and 30 mol% Khafagy et al. (2008), it mostly denotes the variation of different compounds. However, in some articles, variables have been used to represent compound names instead of their values. One such example is REYAlCo where RE = Ce, Pr, Nd, Sm, Gd, Tb, Er, Sc Zhao et al. (2009). This scenario is observed in 1% of the tables.
@@ 3.1.2 Extracting compositions from text
Now, we report the challenges in extracting the compositions reported in the text of the MatSci research papers. We report our statistical findings based on each challenge’s occurrence frequency.
@@
a. Different formats of compositions: The compositions in materials literature does not adhere to a predetermined pattern and encompasses various variations. This is in strict contrast to notations in chemistry where IUPAC nomenclature is used. Some notable examples are as follows.
@@
1. "Erbium-doped glasses with the molar composition 40GeO.10SiO.25NbO.25KO, plus 0.1 to 4 mol% of ErO1.5, were prepared using mixtures of the respective oxides (99.99% purity), with exception of KO, which was added in the form of KCO" Santos et al. (2010).
@@
2. "Bulk samples of (SeTe)Ag (0 x 4) system were prepared by conventional melt quenching technique. High-purity (99.999%) elements with appropriate atomic percentages were sealed in a quartz ampoule (length 100 mm and internal diameter 6 mm) in a vacuum of 10 - 5 mbar" Singh et al. (2012).
@@
3. "The samples having chemical composition of 2(Ca,Sr,Ba)O–TiO–2SiO were examined. CaO, SrO and BaO contents in the samples were varied as shown in Table 1. RO% shows the molar percentage of CaO, SrO or BaO in total RO of CaO+SrO+BaO" Tsuzuku et al. (2002).
@@
b. Extracting variable values in text: Extracting values from variables is challenging since the variable values are specified in different formats with some present only in the text. For instance, consider the following sentence from a peer-reviewed manuscript. Monteiro et al. (2013) A series of tellurite glasses with nominal composition (80-x)TeO–xGeO–10NbO–10KO, where x = 0, 10, 20, 30, 40, 50, 60, 70 and 80 mol%, hereafter named 8T0G, 7T1G, 6T2G, 5T3G, 4T4G, 3T5G, 2T6G, 1T7G and 0T8G, respectively, were doped with 0.2 to 4 mol% .
@@
Here, the values representing the compositions and the respective variable names are present only in the text. A few instances of other formats of composition with variables are shown in Appendix 9.2.2(c).
@@
c. Low recall in extracting compositions expressed with variables: 28% of the articles have compositions written with variables of which 28.57% does not provide any values for the variables in the text. Among the 71.53% where values are present, 40% of them do not mention the step size for the range of values taken by the variable. For example, consider the text representing a set of compositions as follows from a manuscript: x(0.75AgI:0.25AgCl):(1-x)(AgO:WO), where 0.1x1 in molar weight fraction. Agrawal et al. (2004) The step size of 0.1 is mentioned nowhere in the text but could be inferred from the composition table present in the paper. Therefore, extracting only from the text in such cases leads to more errors, and this can be resolved by connecting the variables to the correct composition table containing the variable.
@@
"Lithium disilicate glass was prepared in 30 g quantity by heating stoichiometric homogeneous mixtures of lithium carbonate (99.0%), Synth, and silica (99.9999%), Santa Rosa, for 4 h at 1500 °C in a platinum crucible." Schröder et al. (2014). This text has the composition of the glass as lithium disilicate mentioned indirectly without clearly mentioning the percentages or numbers.
@@
e. Unstable and irrelevant composition extraction: Unstable reagents and other irrelevant compositions which does not refer to the material are also identified as compositions due to a lack of robust parsers. is an unstable entity referring to the aluminium tetrahedral structure, while can be a composition. These undesired extractions can lead to a huge drop in the precision of the IE model, and separating them from the material composition is not an easy task. Only a domain expert with the help of the source article can confirm whether the extraction is relevant or not.
3.1.3 Extracting compositions from table and text jointly:

### v3

@@ Abstract
The discovery of new materials has a documented history of propelling human progress for centuries and more. The behaviour of a material is a function of its composition, structure, and properties, which further depend on its processing and testing conditions. Recent developments in deep learning and natural language processing have enabled information extraction at scale from published literature such as peer-reviewed publications, books, and patents. However, this information is spread in multiple formats, such as tables, text, and images, and with little or no uniformity in reporting style giving rise to several machine learning challenges. Here, we discuss, quantify, and document these challenges in automated information extraction (IE) from materials science literature towards the creation of a large materials science knowledge base. Specifically, we focus on IE from text and tables and outline several challenges with examples. We hope the present work inspires researchers to address the challenges in a coherent fashion, providing a fillip to IE towards developing a materials knowledge base.
@@ 1 Introduction
Understanding a material’s behavior requires knowledge about its composition, properties, processing and testing protocols, and microstructure—represented as the materials science (MatSci) tetrahedron (see Fig. 1). These different aspects of a material are reported by researchers in peer-reviewed publications, patents, and other scientific documents. Recently, there have been several attempts to exploit the advances in machine learning (ML) and artificial intelligence (AI) towards automated information extraction (IE) from literature[1, 2, 3, 4]. These include the development of materials specific language models [5, 6, 7, 8], rule-based systems [9, 10, 11, 12, 13], IE from tables [8, 14, 15], and IE from images [16, 17, 18, 19]. The widely varying information expression styles in research papers makes the automated MatSci IE a challenging task. Most of the works have focused on IE in a specific domain; hence, the transferability to different materials is not explored. Moreover, no consolidated work exists that explores the specific challenges associated with IE in MatSci and the gain associated with solving these challenges, which provides a clear direction to the researchers regarding the areas that require increased attention.
@@
We thoroughly review MatSci articles to identify IE challenges towards completing the materials tetrahedron (see Fig. 1). We also highlight some of the major challenges toward the development of a “universal” MatSci knowledge base linking the extracted information from multiple sources and forms of data—structured, semi-structured, and unstructured. Indeed, millions of scientific documents exist reporting information about various materials known to humans. Thus, the automated development of MatSci IE will lead to a rich knowledge base on materials. The outline of the paper is as follows: First, we explain the methodology of collecting papers for review and annotation process. Then, in the results and discussion sections, we investigate the proportion of each of the entities, such as composition, structure, properties, processing, and testing conditions, reported in tables or text of the articles, followed by the challenges faced in their extraction. We quantify how frequently a challenge occurs to motivate researchers to gauge the amount of information that will be obtained after solving the respective challenges. We further identify the challenges in extracting and connecting the information from text and tables and among different tables belonging to the same MatSci research papers. Note that the challenges reported for extracting compositions from tables are verified by the present IE models, and only those that are unaddressed or solved unsatisfactorily are reported in the main text, whereas some of the existing challenges that have been resolved satisfactorily are documented in the appendix. In our study, DiSCoMaT [8] was employed as the IE model for extracting compositions from tables, recognized as the most effective IE model for this purpose [4]. Concurrently, GPT-4 was utilized to extract compositions from textual content in our study. For extracting properties from MatSci tables, we could not find any domain-specific IE model, but we believe that the challenges reported are valid for any IE models. We have also provided reasons and examples to elaborate on the same. Regarding IE from text to complete materials tetrahedron, we have highlighted examples where existing IE models also make mistakes. Finally, we provide some guidelines for presenting machine and human-friendly tables that enable automated MatSci IE from research papers.
@@ 2 Methodology
To elucidate the challenges, we referred to a dataset of 2536 peer-reviewed publications on MatSci. This dataset is taken from recent work on IE from tables [8], where the authors used distant supervision to annotate tables from research papers based on respective compositions present in INTERGLAD [20]. The tables in val and test data were annotated manually by indicating the relevant rows and columns that should be used to extract material compositions. Fig. 1 shows different sections of the paper where these different components are majorly reported. The statistics of each challenge were computed by randomly taking 50/100 tables from the manually annotated val and test dataset. In cases where this was not applicable, we further performed manual annotation on an additional 50 papers or 100 relevant tables selected randomly from the corpus. For instance, we randomly selected 100 composition tables from the manual annotation in the existing dataset for composition extraction. However, no such manual annotation was available for properties. For this problem, we selected 100 random property tables from the corpus and manually annotated the frequency of the challenges in property extraction. Note that all the challenges and their reported frequencies are based on manual annotation, which is more reliable than any ML-based technique, such as distant supervision. Further, we manually analyzed tables or text for the occurrence of each of the entities, such as composition, structure, and property. All the results and data associated with the annotation process are shared in the following link.
@@ 3.1.1 Extracting compositions from tables
Here, we summarize the major challenges in composition extraction from tables. To this extent, we investigated 100 randomly selected composition tables from the manually annotated data to report the frequency of occurrence of each challenge.
@@
a. Variation in table structure and information content: An analysis of 100 random MatSci composition tables revealed that these tables do not follow any standard structure. Accordingly, following earlier schema proposed by Gupta et al. [8], composition tables can categorized into two broad categories—multi-cell composition (MCC) and single-cell composition (SCC). These are further subdivided into tables containing complete information (CI) and partial information (PI). When the entire composition is written inside a single cell, it is classified as an SCC table, whereas when the composition is written across multiple cells of the table by reporting the value of each constituent (compounds or elements) of the composition in separate cells, it is defined to be an MCC table. If the table contains all the information regarding the constituents of the material, they are classified as CI tables (complete information). Alternatively, if only some of the constituents are mentioned in the table for the material, they are PI tables. In the latter case, we need to extend the analysis to the text of the article to extract the full composition. Fig. 3 illustrates all 4 types of tables [21, 22, 23, 24]. The most prevalent composition table types are MCC-CI (36%), followed by SCC-CI (30%). PI tables are less common, with 24% being MCC-PI and the remaining 10% being SCC-PI. Note that this distribution may also vary significantly depending on the material types. For instance, it is common practice in alloys to skip the major element while describing the composition in a table. In previous work by Gupta et al. [8], while an F1 score of 78.21% and 65.41% have been achieve for extraction from SCC-CI and MCC-CI tables, respectively, an F1 score of only 51.66% has been achieved for extraction from MCC-PI. Although the researchers have not explicitly focused on SCC-PI, we used their best model for SCC-PI tables and obtained 47.19% as the F1 score. Hence, there is a significant scope for improvement in extracting compositions from PI tables.
@@
b. Presence of nominal and experimental compositions: While the nominal composition is the amount of chemicals taken initially to prepare the material, analyzed/experimental composition refers to the actual composition of material obtained after analyzing the manufactured material (see Fig. 4(a)) [25, 26]. Our analysis revealed that in 3% of the tables, both nominal and analyzed/experimental compositions are reported. These values are not reported in any fixed pattern, making it difficult to correctly separate the nominal and analyzed compositions after extraction.
@@
c. Compositions and related info inferred from other documents: In some tables, the details of the glasses studied are not explicitly mentioned; rather, references to previous research publications which use the same material are provided in the tables or their captions (see Fig. 4(b)). Thus, the composition or the other associated information of the material which is missing in the current publication must be extracted from the cited work, which then must be combined with the relevant information of the material present in the current work. We found references about different entities of the material in 11 tables [27, 28]. 4 out of the 11 tables have not explicitly mentioned compositions, due to which the IE model [8] was unsuccessful in obtaining the desired compositions.
d. Composition inferred from material ID: We observed that 10% of the total composition tables contain IDs with essential material composition information. In 60% of these tables, DiSCoMaT [8] failed to extract the compositions correctly. Most of these tables did not mention the materials’ composition separately, thereby making the extraction challenging. For example, some of the materials have their compositions indicated within the IDs in an abbreviated form [29] and did not mention them explicitly (see Fig. 5(a)). We also found tables where the composition of the materials is not specified; instead, their standard names are used as IDs. Such examples include Wollastonite or Diopside [30], which have a fixed chemical composition that can be obtained from standard sources/databases. In some cases, the composition was specified separately, but the IE model failed to extract the composition correctly due to dependency on material IDs, as shown in 5(b). Here, the variable ‘M’ needs to be substituted by elements like ‘W’, ‘Nb’, or ‘Pb’, which needs to be inferred using the material IDs mentioned in the first column of the illustrated table [31].
@@
e. Variables used to represent compounds: When a composition is expressed with variables such as TeO2+15B2O3+15P2O5+Li2O, where = 5, 10, 15, 20, 25 and 30 mol% [32], it mostly denotes the variation of different compounds. However, in some articles, variables have been used to represent compound names instead of their values. One such example is RE36Y20Al24Co20 where RE = Ce, Pr, Nd, Sm, Gd, Tb, Er, Sc [33]. This scenario is observed in 1% of the tables, where DiSCoMaT [8] fails to extract the material compositions. Note that this particular case can be solved using GPT-4, but as DiSCoMaT performs better in composition extraction from tables than GPT-4 [4], and a pipeline of GPT-4 and DiSCoMaT is not feasible, hence, this still remains an open challenge.
@@ 3.1.2 Extracting compositions from text
Now, we discuss the challenges in extracting the compositions reported in the text of the MatSci research papers. We report our statistical findings based on the frequency of each challenge. We also use GPT-4 to extract the compositions from text. The prompts given to GPT-4 for composition extraction are provided in Table LABEL:tab:gpt4. Specifically, we have used gpt4-1106 model through the OpenAI Python library. The temperature was set to 0.0 for reproducibility.
@@
a. Different formats of compositions: The compositions in materials literature do not adhere to a predetermined pattern and encompass several variations. This is in strict contrast to notations in chemistry, where IUPAC nomenclature is used. Some notable examples are as follows.
@@
1. "Erbium-doped glasses with the molar composition 40GeO2.10SiO2.25Nb2O5.25K2O, plus 0.1 to 4 mol% of ErO1.5, were prepared using mixtures of the respective oxides (99.99% purity), with exception of K2O, which was added in the form of K2CO3" [34].
@@
2. "Bulk samples of (Se80Te20)100-xAgx (0 x 4) system were prepared by conventional melt quenching technique. High-purity (99.999%) elements with appropriate atomic percentages were sealed in a quartz ampoule (length 100 mm and internal diameter 6 mm) in a vacuum of 10 - 5 mbar" [35].
@@
3. "The samples having chemical composition of 2(Ca,Sr,Ba)O–TiO2–2SiO2 were examined. CaO, SrO, and BaO contents in the samples were varied as shown in Table 1. RO% shows the molar percentage of CaO, SrO or BaO in total RO of CaO+SrO+BaO" [36].
@@
b. Extracting variable values in text: Extracting values from variables is challenging since the variable values are specified in different formats, with some present only in the text. For instance, consider the following sentence from a peer-reviewed manuscript. [37] A series of tellurite glasses with nominal composition (80-x)TeO2–xGeO2–10Nb2O5–10K2O, where x = 0, 10, 20, 30, 40, 50, 60, 70 and 80 mol%, hereafter named 8T0G, 7T1G, 6T2G, 5T3G, 4T4G, 3T5G, 2T6G, 1T7G and 0T8G, respectively, were doped with 0.2 to 4 mol% .
Although GPT-4 understands the doping element, since the entire information is not present in the same sentence and the exact values of doping content are not specified, it does not extract the composition successfully.
@@
Here, the values representing the compositions and the respective variable names are present only in the text. Appendix 9.2.2(c) shows a few instances of other composition formats with variables. However, it may be noted that if full information is present in the sentences, GPT-4 is able to extract information correctly for the cases where the compositions are given in the form of variables.
@@
c. Low recall in extracting compositions expressed with variables: 28% of the articles have compositions written with variables, of which 28.57% does not provide any values for the variables in the text. Among the 71.53% where values are present, 40% of them do not mention the step size for the range of values taken by the variable. For example, consider the text representing a set of compositions as follows from a manuscript: x(0.75AgI:0.25AgCl):(1-x)(Ag2O:WO3), where 0.1x1 in molar weight fraction. [38] The step size of 0.1 is mentioned nowhere in the text but could be inferred from the composition table present in the paper. Therefore, extracting only from the text in such cases leads to more errors, and this can be resolved by connecting the variables to the correct composition table containing the variable. GPT-4 takes the endpoints for substituting the values in the compositions. However, due to a lack of information, it does not extract complete compositions due to the lack of values between the extreme values.
@@
"Lithium disilicate glass was prepared in 30 g quantity by heating stoichiometric homogeneous mixtures of lithium carbonate (99.0%), Synth, and silica (99.9999%), Santa Rosa, for 4 h at 1500 °C in a platinum crucible." [39]. This text indirectly mentions the glass’s composition as lithium disilicate without clearly mentioning the percentages or numbers. GPT-4 is able to infer the chemical formulas from chemical names but cannot infer the exact composition and its percentages from the sentence.
@@
e. Unstable and irrelevant composition extraction: Unstable reagents and other irrelevant compositions which does not refer to the material are also identified as compositions due to a lack of robust parsers. is an unstable entity referring to the aluminum tetrahedral structure, while can be a composition. These undesired extractions can lead to a huge drop in the precision of the IE model, and separating them from the material composition is not easy. Only a domain expert, with the help of the source article, can confirm whether the extraction is relevant or not.
GPT4 fails to differentiate compositions from unstable compounds.
It is worth noting that although GPT-4 can address some of these challenges, especially extraction from text, its closed nature makes it challenging to use it at scale and for custom applications. Some of the reasons are:
1. Often, the research documents could be highly sensitive, preventing their sharing with commercial models such as GPT-4.
2. The inability of GPT-4 to be combined with smaller predictive models like DiSCoMaT prevents exploiting excellent domain-specific models that extract information very accurately.
3.
The commercial nature of such models can make it prohibitive due to the expenses associated with the usage due to the large number of sentences to be analyzed in the research papers and any additional prompt-engineering involved.
Therefore, GPT-4 may not be an ideal baseline for IE at large scale from research publications.
3.1.3 Extracting compositions from table and text jointly